The percent error in a measurement is the ratio of the error in the measurement to the measurement​ itself, expressed as a percent. When writing a computer​ program, the memory remaining is determined as 2650 bytes and then it is correctly found to be  2700 bytes. What is the percent error in the first​ reading?

Answer by Theo(13342)

the first reading is 2650
the correct reading is 2700.
the error is 50.
the percent error is 50/2700 = .0185185185 * 100 = 1.85185185 percent of the correct reading.

if you took a ratio of the first reading to the correct reading, you would get 2650 / 2700 = .9814814815 which is equal to 98.14814815 percent.

the first reading is equal to 98.14814815 percent of the correct reading.

since the first reading is 98.14814815 percent of the correct reading and since the error is 1.85185185 percent of the correct reading below it, if you add the two of these together, you should get 100% of the correct reading.

1.85185185 percent of the correct reading + 98.14814815 percent of the correct reading = 100 percent of the correct reading.

i believe that your solution is that the error is 1.85185185 percent of the correct reading.